At $11.25 per foot for the curved portion and $1.75 per foot for the straight portion. How much will it cost to put molding around a window. Use 3.14
Answer by checkley71(8403)

4 FEET 6 INCHES=4.5 FEET.
2 FEET 3 INCHES=2.25 FEET.
4.5*2=9 FEET FOR THE STRAIGHT PORTION.
9*1.75=$15.75 IS THE COST FOR THE STRAIGHT MOLDING.
RADIUS OF THE TOP SECTION IS 2.25 FEET.
1/2 THE CIRCUMFERANCE IS THE LENGTH OF THE CURVED SECTION FOR MOLDING.
C=2*PI*R
C=2*3.14*2.25
C=14.13 IS THE CIRCUMFERANCE.
1/2*14.13=7.065 FEET IS HALF THE CIRCUMFERANCE.
7.065*11.25=79.48 COST OF THE CURVED SECTION.
15.75+79.48=$95.23 FOR THE TOTAL COST.
DON'T KNOW HOW THEY GOT $103.11?